The pvscript package^{∗} for use with L^{A}TEX2e

Gerd Neugebauer Mainzer Str. 16 56321 Rhens

             gerd@informatik.uni-koblenz.de

1996/07/18

1 Introduction

The font pvscript provides a handwritten font. It can be found on the CTAN in the directory tex-archive/fonts/script. This package provides means to use this font.

2 Usage

This file can be used as a package by placing its name in the argument of \usepackage. Afterwards the font family pvscript is defined. This could also have been done by providing a font definition file.

\pvscript The command \pvscript changes the current font family to pvscript and the encoding to OT1. Usually this should be used in a TEX group only since it changes the font encoding and redefines the umlaut macro.

\textpvscript The macro \textpvscript typesets its argument in the pvscript font.

3 Implementation

First we declare a new font family for the pvscript font.

 1 \DeclareFontFamily{OT1}{pvscript}{}

This font is available in the normal shape and in a bold variant. Here we can get the desired font by (silently) scaling the present fonts.

 2 \DeclareFontShape{OT1}{pvscript}{m}{n}{<-> s * [12.84]script10}{}
 3 \DeclareFontShape{OT1}{pvscript}{bx}{n}{<-> s * [12.84]scriptb10}{}
 4 \DeclareFontShape{OT1}{pvscript}{b}{n}{<-> sub * pvscript/bx/n}{}

Next we define some macros to perform the mapping of umlaut characters which have to be combined out of two characters each.

 5 \newcommand\pvscript@a{\mbox{\rlap{\kern.25em\symbol{34}}a}}

∗This file has version number v1.2, dated 1996/07/18.

 6 \newcommand\pvscript@o{\mbox{\rlap{\kern.1em\symbol{34}}o}}
 7 \newcommand\pvscript@u{\mbox{\rlap{\kern.1em\symbol{34}}u}}
 8 \newcommand\pvscript@A{\mbox{\rlap{\kern.5em\raisebox{1ex}{\symbol{34}}}A}}
 9 \newcommand\pvscript@O{\mbox{\rlap{\kern.4em\raisebox{1ex}{\symbol{34}}}O}}
10 \newcommand\pvscript@U{\mbox{\rlap{\kern.4em\raisebox{1ex}{\symbol{34}}}U}}

Now we define the font changing command.

pvscript The macro \pvscript selects the pvscript family.

11 \DeclareRobustCommand\pvscript{
12 \let\pvscript@original@umlaut=\"
13 \def\"##1{%
14  \@ifundefined{pvscript@##1}{%
15  \pvscript@original@umlaut{##1}}{%
16  \@nameuse{pvscript@##1}}}%
17 \def\ss{ss}%
18 \fontfamily{pvscript}%
19 \fontencoding{OT1}%
20 \selectfont}

textpvscript The macro \textpvscript typesets its argument in the pvscript font.

21 \newcommand\textpvscript[1]{\begingroup\pvscript #1\endgroup}